### v2

Abstract
Two decades of research on the euro’s trade effects have produced estimates ranging from 4% to 30%, with no consensus on the magnitude. We find evidence that this divergence may reflect genuine heterogeneity in the euro’s trade effect across country pairs rather than methodological differences alone. Using Eurostat data on 15 EU countries (12 eurozone members plus Denmark, Sweden, and the UK as controls) from 1995–2015, we estimate that euro adoption increased bilateral trade by 29% on average (14.1% after fixed effects correction), but effects range from $12\%$ to $79\%$ across eurozone pairs. Core eurozone pairs (e.g., Germany–France, Germany–Netherlands) show large gains, while peripheral pairs involving Finland, Greece, and Portugal saw smaller or negative effects, with some negative estimates statistically significant and interpretable as trade diversion. Pre-euro trade intensity and GDP account for over 90% of feature importance in explaining this heterogeneity. Extending to EU28, we find evidence that crisis-era adopters (Slovakia, Estonia, Latvia) pull down naive estimates to 4.3%, but accounting for fixed effects recovers estimates of 13.4%, consistent with the EU15 fixed-effects baseline of 14.1%. Illustrative counterfactual analysis suggests non-eurozone members would have experienced varied effects: UK (+33%), Sweden (+22%), Denmark (+19%). The wide range of prior estimates appears to be largely a feature of the data, not a bug in the methods.
keywords:
euro , currency union , trade , heterogeneous treatment effects , gravity model , synthetic control
1 Introduction
Two decades of research on the euro’s trade effects have failed to reach consensus. Gravity model estimates cluster around 4–6%, while synthetic control studies find effects as high as 30% for specific country pairs. Rose (2016) shows that larger datasets produce systematically larger estimates. Gunnella et al. (2021) attribute the divergence to methodological differences. We propose an alternative explanation: the wide range may reflect genuine heterogeneity in treatment effects across country pairs, which existing methods cannot capture.
Understanding this heterogeneity has direct policy relevance. Sweden, Denmark, and the United Kingdom opted out of the eurozone, and policymakers in these countries have periodically revisited the adoption question. If the euro’s trade effects are uniformly positive, the trade-based case for adoption is straightforward. But if effects vary substantially—large for some pairs, near-zero for others—the calculus becomes more complex. The same logic applies to prospective members in Central and Eastern Europe. Knowing which country pairs benefit most, and why, is essential for informed policy.
We estimate the full distribution of euro trade effects across all eurozone country pairs. Our approach suggests substantial heterogeneity that may help reconcile the divergent estimates in prior literature: gravity models, which estimate average effects, capture the center of the distribution (around 20%), while synthetic control studies, which focus on specific high-trade pairs, capture the upper tail (30%+)—precisely because SCM studies typically select pairs where credible synthetic counterfactuals exist. We also generate illustrative counterfactual predictions for non-eurozone EU members, estimating what trade effects Sweden, Denmark, and the United Kingdom might have experienced had they adopted the euro.
The EU provides a useful setting for this analysis because member countries meet the Maastricht criteria for fiscal discipline, price stability, and exchange rate stability, yet not all have adopted the euro. This creates a comparison where the main observable difference between treatment and control groups is euro adoption itself, though we acknowledge that unobserved differences may remain. To our knowledge, this is among the first studies to estimate the full distribution of heterogeneous treatment effects of currency union membership. Methodologically, we use causal forests with double machine learning to estimate conditional average treatment effects at the country-pair level.
The paper proceeds as follows. Section 2 reviews the literature on currency unions and trade. Section 3 describes the causal forest methodology. Section 4 presents the data. Section 5 reports results on heterogeneous treatment effects. Section 6 discusses the findings. Section 7 presents counterfactual analysis for non-eurozone EU members. Section 8 concludes.
2 Literature Review
Research on the trade effects of common currencies began with the influential contribution of Rose (2000), which used a gravity framework and reported very large trade increases among currency-union members. This result triggered extensive debate about identification, selection into currency unions, and the credibility of cross-sectional comparisons. Addressing these concerns, Persson (2001) applied matching-based reasoning and argued that differences between treated and untreated pairs can generate substantial upward bias in naïve estimates.
A second wave of work emphasized panel variation and dynamics. For example, Bun and Klaassen (2002) estimated a dynamic panel model focused on the euro and found a modest short-run effect that accumulates over time. In parallel, methodological advances in gravity estimation improved the interpretability of currency-union coefficients. Anderson and van Wincoop (2003) formalized multilateral resistance—showing that bilateral trade depends on relative trade costs—and their framework became a benchmark for modern gravity specifications.
Direct evidence on the euro specifically emerged early. Micco et al. (2003) provided early EMU evidence using developed-country data and found positive trade effects. Subsequent gravity work increasingly adopted estimation strategies designed to handle heteroskedasticity and zero trade flows. Santos Silva and Tenreyro (2006) showed that log-linear OLS can be misleading under heteroskedasticity and advocated PPML-type approaches that have since become standard in gravity estimation.
Despite improvements in data and methods, estimates of the euro’s trade effect continued to vary. One explanation is that results are sensitive to data coverage and specification choices: Rose (2016) documented that larger datasets (more countries/longer spans) tend to deliver systematically larger EMU trade estimates. A complementary explanation emphasizes that different empirical approaches target different parameters. For example, Gunnella et al. (2021) compare gravity and synthetic control approaches and report modest “average-type” gravity effects alongside substantially larger effects for some pairs under synthetic control.
A major methodological branch in this literature uses Synthetic Control Methods (SCM), which construct a data-driven counterfactual from weighted donor units and are particularly attractive when a small number of treated units receive a discrete policy intervention. The approach is grounded in foundational contributions such as Abadie et al. (2010). SCM has been applied both within and beyond trade settings to study institutional and macro-financial regime changes. Aytuğ (2017) uses SCM to construct counterfactual exchange rate volatility in the absence of Turkey’s Reserve Option Mechanism, illustrating that policy effects can be strongly state-dependent and sensitive to concurrent monetary tightening. In trade-policy evaluation, Aytuğ et al. (2017) apply SCM to assess the EU–Turkey Customs Union and show sizable effects relative to a synthetic counterfactual. In the euro context, Saia (2017) uses SCM to study the counterfactual of UK euro adoption and finds sizable trade gains in that scenario.
At the same time, the SCM toolkit has expanded in ways that matter for interpretation. Ben-Michael et al. (2021) propose the Augmented Synthetic Control Method to improve performance when perfect pre-treatment fit is not feasible. Arkhangelsky et al. (2021) introduce Synthetic Difference-in-Differences, combining elements of DID and synthetic control to improve robustness in common empirical settings. More recently, Di Stefano and Mellace (2024) propose an “inclusive” synthetic control variant designed for settings where spillovers or indirect effects may contaminate donor pools. These developments strengthen the credibility of comparative-case approaches, but they also reinforce a core limitation: SCM-style designs remain most naturally suited to case-specific evaluation (one or a few treated units) rather than systematic characterization of effects across the full set of euro-area country pairs.
This brings the literature to a central unresolved issue: heterogeneity. Gravity models typically estimate an average (or average-like) effect of euro membership across treated pairs, while SCM studies often focus on a subset of country-pairs where credible synthetic counterfactuals can be built—precisely the setting in which large effects are more likely to be detected. The combination of these approaches therefore leaves open an important question: what is the full distribution of the euro’s trade effects across all euro-area pairs, and which observable features predict where a given pair lies in that distribution?
To address this gap, we build on the growing literature that applies modern causal inference tools to uncover treatment effect heterogeneity. Athey and Imbens (2016) develop tree-based approaches designed specifically for heterogeneous causal effects, providing a foundation for forest-based estimators. Related work on machine-learning-based causal inference, such as Chernozhukov et al. (2018), provides tools for valid inference on causal parameters in high-dimensional settings. Our contribution is to use these ideas to estimate the distribution of euro trade effects across country pairs, thereby reconciling why different methods in the existing literature can yield apparently conflicting results.
Table 1 summarizes the key methodological differences between existing approaches and our causal forest method.
Table 1: Comparison of Estimation Approaches
Gravity/PSM SCM Causal Forest
Estimates Single ATE Few pair effects Full CATE distribution
Heterogeneity Pre-specified only Not systematic Data-driven discovery
Selection bias PSM addresses Matching-based DML residualization
Inference Clustered SE Placebo tests Forest-based CI
Literature range 4–30% 16–30%
3 Methodology
3.1 Existing Approaches: Gravity and Synthetic Control
The gravity model is the workhorse of empirical trade analysis. Following Anderson and van Wincoop (2003) and best practices outlined in Gunnella et al. (2021), the structural gravity equation with a saturated set of fixed effects is:
$\ln(X_{ijt})=\lambda_{it}+\psi_{jt}+\mu_{ij}+\boldsymbol{\beta}^{\prime}\mathbf{z}_{ijt}+\gamma\text{CU}_{ijt}+\varepsilon_{ijt}$ (1)
where $X_{ijt}$ denotes exports from country $i$ to country $j$ at time $t$, $\lambda_{it}$ and $\psi_{jt}$ are exporter-time and importer-time fixed effects that control for multilateral resistance, $\mu_{ij}$ are pair fixed effects that absorb time-invariant bilateral factors (distance, common language, contiguity), $\mathbf{z}_{ijt}$ includes time-varying bilateral controls such as regional trade agreements, and $\text{CU}_{ijt}$ is a dummy equal to one if both countries are in a currency union at time $t$. The coefficient $\gamma$ captures the average treatment effect of euro adoption. Estimation via PPML addresses heteroskedasticity and zero trade flows (Santos Silva and Tenreyro, 2006).
The synthetic control method (SCM) offers an alternative approach by constructing a counterfactual for each treated unit as a weighted combination of control units. Following Abadie and Gardeazabal (2003), the causal effect is estimated as:
$\hat{\tau}_{1}=\mathbf{y}_{1}-\mathbf{y}_{0}^{*}=\mathbf{y}_{1}-\mathbf{Y}_{0}\mathbf{w}^{*}$ (2)
where $\mathbf{y}_{1}$ is the outcome vector for the treated unit, $\mathbf{Y}_{0}$ is a matrix of outcomes for $J$ control units, and $\mathbf{w}^{*}$ is a vector of weights chosen to minimize the distance between treated and synthetic control units in the pre-treatment period. Gunnella et al. (2021) apply SCM to estimate euro effects for specific country pairs, finding effects around 30%. Saia (2017) uses SCM to estimate what UK trade would have been under euro adoption, finding a 16% effect. While SCM provides credible counterfactuals for individual cases, it does not scale to estimate effects for all country pairs or systematically explore heterogeneity.
Both approaches have limitations for our research question. Gravity models estimate a single average effect, potentially masking substantial heterogeneity across pairs. SCM can estimate pair-specific effects but requires selecting cases ex ante and does not identify what drives variation. We propose causal forests as a method that combines the strengths of both: estimating the full distribution of treatment effects while identifying the characteristics that explain heterogeneity.
3.2 From PSM to Causal Forests
Traditional propensity score matching (PSM) addresses selection bias by matching treated and control units based on their probability of treatment. Following Persson (2001), the propensity score is defined as:
$e(X)=P(\text{Euro}_{ijt}=1\mid X_{ijt})$ (3)
where $X_{ijt}$ includes GDP and other pair characteristics. The average treatment effect on the treated (ATT) is then:
$\tau_{\text{ATT}}=E[Y^{(1)}-Y^{(0)}\mid\text{Euro}=1]$ (4)
While PSM addresses selection bias, it estimates only a single average effect. Causal forests extend this framework by estimating heterogeneous treatment effects — allowing the effect to vary with observed characteristics:
$\tau(x)=E[Y^{(1)}-Y^{(0)}\mid X=x]$ (5)
This Conditional Average Treatment Effect (CATE) captures how the euro’s impact varies across country pairs with different characteristics.
3.3 Intuition: How Causal Forests Discover Heterogeneity
Before presenting the technical details, we provide intuition for how causal forests estimate heterogeneous treatment effects. The key insight is that causal forests are designed to answer a different question than traditional methods: rather than asking “what is the average effect?” they ask “for which units is the effect largest, and why?”
A standard decision tree predicts outcomes by recursively splitting the data into groups that are most similar in their outcome values. A causal tree instead splits the data into groups that are most different in their treatment effects. Consider a simple example: if the euro’s effect on trade is larger for country pairs with high pre-existing trade intensity, the causal tree will split on trade intensity, creating one group (high-trade pairs) with a high estimated effect and another group (low-trade pairs) with a lower effect.
The algorithm proceeds as follows. At each node, the tree considers all possible splits of the data (e.g., GDP above/below median, pre-euro trade high/low) and selects the split that maximizes the difference in treatment effects between the resulting subgroups. This process continues recursively until the subgroups become too small. The treatment effect for any unit is then estimated as the average effect within its terminal node (leaf).
A single tree would be noisy and sensitive to the particular sample. Causal forests address this by growing many trees (we use 500) on bootstrapped samples and averaging their predictions. This ensemble approach reduces variance while preserving the ability to capture complex heterogeneity patterns.
Two features distinguish causal forests from standard machine learning methods. First, “honest” estimation uses separate subsamples for determining the tree structure and estimating effects within leaves, preventing overfitting and enabling valid statistical inference. Second, the algorithm provides not just point estimates but confidence intervals for each unit’s treatment effect, allowing researchers to assess statistical significance at the individual level.
For our application, this means the causal forest can discover — without prior specification — that the euro’s trade effect is large for core European pairs with deep existing trade relationships, modest for peripheral pairs, and near-zero for pairs involving Greece. Traditional methods would require the researcher to hypothesize these patterns in advance and test them through subgroup analysis or interaction terms. Figure 1 illustrates this process.
[…]
Figure 1: Illustration of causal tree splitting. The tree recursively partitions country pairs into subgroups with different treatment effects. At each node, the split is chosen to maximize the difference in euro effects between subgroups. Terminal nodes (leaves) show estimated effects ranging from +77% for high-trade, high-GDP pairs (e.g., France–Luxembourg) to $3\%$ for low-trade, low-GDP pairs (e.g., Finland–Ireland). Colors indicate effect magnitude: green (high), yellow (medium), red (low). A causal forest averages predictions across many such trees.
3.4 Causal Forest with Double Machine Learning
We apply Causal Forests with Double Machine Learning (Athey and Wager, 2018; Chernozhukov et al., 2018) to estimate the full distribution of CATEs. The CausalForestDML estimator proceeds in three steps:
Step 1: Nuisance estimation. Use machine learning to estimate:
$\hat{m}(W)$ $=E[Y\mid W]\quad\text{(outcome model)}$ (6)
$\hat{e}(W)$ $=E[T\mid W]\quad\text{(propensity score)}$ (7)
Step 2: Residualize. Partial out the effect of controls:
$\tilde{Y}$ $=Y-\hat{m}(W)$ (8)
$\tilde{T}$ $=T-\hat{e}(W)$ (9)
Step 3: Causal forest. Estimate $\tau(X)$ using an honest random forest on the residualized data $(\tilde{Y},\tilde{T},X)$.
This approach removes regularization bias from the ML first stage and provides valid confidence intervals. It is doubly robust: consistent if either $\hat{m}$ or $\hat{e}$ is correctly specified. We implement CausalForestDML with Random Forest first-stage models (200 trees, min_samples_leaf = 20), a causal forest with 500 trees and honest splitting (min_samples_leaf = 30), and a Random Forest Classifier for the binary treatment model.
3.5 Constructing Counterfactuals
A fundamental challenge in causal inference is that we never observe the counterfactual outcome — what trade would have been for a eurozone pair had they not adopted the euro, or for a non-eurozone pair had they adopted it. Using the potential outcomes framework, each unit has two potential outcomes: $Y^{(1)}$ (outcome if treated) and $Y^{(0)}$ (outcome if not treated). The treatment effect for unit $i$ is:
$\tau_{i}=Y_{i}^{(1)}-Y_{i}^{(0)}$ (10)
We observe $Y_{i}^{(1)}$ for treated units and $Y_{i}^{(0)}$ for control units, but never both for the same unit.
Traditional matching methods address this by finding “twin” control units with similar observable characteristics to impute the missing potential outcome. Causal forests take a different approach: they estimate the conditional average treatment effect function $\tau(x)=E[Y^{(1)}-Y^{(0)}\mid X=x]$ from the data, then apply this learned function to predict effects for any unit.
For a non-eurozone pair like UK–Germany with characteristics $x_{\text{UK-DE}}$, the counterfactual prediction is:
$\hat{\tau}(\text{UK-DE})=\hat{\tau}(x_{\text{UK-DE}})$ (11)
The causal forest predicts what effect euro adoption would have had based on the pair’s characteristics (GDP, pre-euro trade intensity) and how similar characteristics related to treatment effects among eurozone pairs. The key identifying assumption is that the CATE function learned from treated units generalizes to untreated units with similar characteristics — formally, that $\tau(x)$ is the same for treated and untreated units conditional on $X=x$.
3.6 Treatment Definition and Identification
We define the treatment as both countries in a pair having adopted the euro:
$T_{ijt}=\mathbf{1}[\text{Euro}_{i}=1]\times\mathbf{1}[\text{Euro}_{j}=1]$ (12)
The sample is restricted to pairs where both countries are EU members, ensuring all units faced a credible possibility of treatment. Countries outside the EU (e.g., United States, Japan, Switzerland) could never adopt the euro and thus do not constitute a valid counterfactual. Within the EU sample, eurozone members (Austria, Belgium, Finland, France, Germany, Greece, Ireland, Italy, Luxembourg, Netherlands, Portugal, Spain) constitute the treatment group, while EU members that opted out (Denmark, Sweden, United Kingdom) constitute the control group.
The outcome variable is log real bilateral trade:
$Y_{ijt}=\ln\left(\frac{\text{Exports}_{ij}+\text{Imports}_{ij}}{\text{PPI}_{t}}\right)$ (13)
For PPML specifications in the gravity benchmark, we use trade in levels following Santos Silva and Tenreyro (2006). Effect modifiers $X$ — the variables that may drive heterogeneity in treatment effects — include log GDP product, log GDP per capita, and pre-euro trade intensity (average bilateral trade 1995–1998, the pre-treatment period in our sample).
Controls $W$ used in the first-stage nuisance estimation address potential confounders that affect both euro adoption and trade:
  • 1.
    Log GDP product: Larger economies trade more and were more likely to be founding eurozone members.
  • 2.
    Log GDP per capita: Richer countries have different trade patterns and faced different incentives for euro adoption.
  • 3.
    Year: Controls for business cycle effects and common time trends; the DML residualization removes year-specific variation from both outcome and treatment, analogous to year fixed effects.
Identification relies on the conditional independence assumption: conditional on the controls $W$, treatment assignment is independent of potential outcomes. The DML framework addresses selection on observables by flexibly modeling the relationship between confounders and both the outcome and treatment. The causal forest then estimates heterogeneous effects conditional on effect modifiers $X$, which may overlap with but are conceptually distinct from the confounders $W$.
3.7 Threats to Identification
While the DML framework addresses selection on observables through flexible first-stage estimation, our identification strategy faces potential threats from unobserved confounders. We discuss the main concerns and the evidence bearing on them.
Countries that adopted the euro may have had stronger political commitment to European integration, which could independently affect trade through non-tariff barrier reductions, regulatory harmonization, or business confidence. However, all countries in our sample are EU members that met the Maastricht criteria, suggesting similar baseline commitment to integration. The non-adopters (UK, Sweden, Denmark) obtained formal opt-outs, reflecting specific domestic political constraints rather than weaker commitment to European trade integration per se. Denmark maintains a currency peg to the euro, demonstrating commitment to monetary stability without formal adoption.
Higher-quality institutions might facilitate both euro adoption and trade expansion. We partially address this by restricting the sample to EU members, which share common institutional frameworks including the single market, common external tariff, and harmonized regulations. The remaining variation in institutional quality within the EU is modest compared to cross-country studies that include developing economies.
Countries experiencing economic booms might have been more likely to adopt the euro and to experience trade growth. We address this through year fixed effects in the DML first stage, which remove common time trends. The EU15 sample benefits from synchronized adoption timing (1999–2001), limiting variation in business cycle position at treatment. The EU28 extension, where adoption timing varies with the 2008–2012 crisis, shows how business cycle confounding can bias estimates downward—and how CFFE correction recovers consistent estimates.
An ideal instrument would predict euro adoption but affect trade only through the adoption channel. Potential instruments face validity concerns:
  • 1.
    Geographic proximity to Brussels: Correlated with trade through standard gravity channels
  • 2.
    Historical currency arrangements: Reflect deep economic integration that directly affects trade
  • 3.
    Referendum outcomes: Endogenous to economic conditions and trade expectations
  • 4.
    Political party composition: May affect trade policy directly through non-monetary channels
Following Barro and Tenreyro (2007), who use the probability of independently adopting a third country’s currency as an instrument, we considered similar approaches but found them inapplicable to the eurozone context where adoption was a coordinated political decision rather than independent currency choices.
We provide two forms of evidence supporting the parallel trends assumption. First, an event study analysis (Figure A.1) shows that pre-treatment coefficients (1995–1997) are small in magnitude ($3\%$ to $5\%$) compared to post-treatment effects ($8\%$ to $24\%$), with a clear break at 1999 when effects become positive. Second, placebo tests (Figure A.2) assigning fake treatment dates (1995, 1997) find no significant “effects,” suggesting our estimates do not reflect pre-existing differential trends.
Following Oster (2019), we assess how much selection on unobservables would be required to explain away our results (Table A.3). The analysis suggests that unobserved confounders would need to be substantially more important than the observed confounders (GDP, GDP per capita, year) to reduce the estimated effect to zero. Given that our controls capture the main economic determinants of both euro adoption and trade, this degree of omitted variable bias appears implausible.
Leave-one-out analysis (Figure A.4) dropping each country in turn shows that no single country drives the results. The ATE remains stable within the confidence interval of the full-sample estimate regardless of which country is excluded, including Luxembourg (which has the largest estimated effects) and peripheral economies like Greece and Portugal.
In summary, while we cannot definitively rule out all unobserved confounding, the combination of (1) restricting to EU members with similar institutional frameworks, (2) flexible DML adjustment for observable confounders, (3) pre-trends evidence supporting parallel trends, (4) placebo tests finding no spurious effects, and (5) sensitivity analysis suggesting implausible degrees of omitted variable bias provides reasonable confidence in our identification strategy.
4 Data
We use bilateral trade data from Eurostat, which provides comprehensive trade statistics for EU member states. The dataset covers 15 EU countries (EU15) from 1995 to 2015, matching the methodology of Gunnella et al. (2021). While Eurostat trade data extends back to 1988, GDP data from Eurostat’s national accounts becomes consistently available only from 1995, making earlier years unsuitable for our analysis which requires GDP controls. We restrict the sample to the EU15 (pre-2004 enlargement members) for two reasons. First, recent euro adopters among the 2004+ enlargement countries have insufficient post-treatment data: Slovenia joined in 2007 (8 years), Slovakia in 2009 (6 years), Estonia in 2011 (4 years), Latvia in 2014 (1 year), and Lithuania in 2015 (0 years in our sample). Second, this sample matches the sample used by Gunnella et al. (2021), enabling direct comparison. The outcome, treatment, and control variables are defined in Section 3.4.
The eurozone members in our sample are Austria, Belgium, Finland, France, Germany, Greece, Ireland, Italy, Luxembourg, Netherlands, Portugal, and Spain. The control group consists of EU members that did not adopt the euro: Denmark, Sweden, and the United Kingdom. This design avoids reliance on non-European donor countries, a common concern in early euro studies that used global samples.
Table 2 reports summary statistics. Log real trade has a mean of 22.30 with substantial variation (standard deviation of 1.73), reflecting heterogeneity in bilateral trade flows across EU pairs. The euro adoption indicator equals one for 51% of observations, reflecting the balanced panel structure with eurozone pairs observed both before and after 1999. Table 3 shows the panel contains 2,149 observations covering 105 unique country pairs across 15 countries. Of these, 1,100 (51.2%) are treated while 1,049 (48.8%) serve as controls.
Table 2: Summary Statistics
Variable Mean Std. Dev. Min Max
Log real trade 22.30 1.73 15.99 26.06
Log GDP product 25.74 1.61 21.37 29.73
Log GDP per capita 20.50 0.64 18.31 22.45
EU membership 1.00 0.00 1.00 1.00
Euro adoption 0.51 0.50 0.00 1.00
Table 3: Sample Composition
Characteristic Value
Total observations 2,149
Unique country pairs 105
Countries 15
Year range 1995–2015
Treated (euro = 1) 1,100 (51.2%)
Control (euro = 0) 1,049 (48.8%)
Mean bilateral trade EUR 16.7B
Median bilateral trade EUR 4.9B
Table 4 compares covariate means between eurozone and non-eurozone pairs in the post-1999 period. Non-eurozone pairs (involving Denmark, Sweden, or UK) have higher GDP per capita (standardized difference of 0.46-0.46) and larger combined GDP (0.23-0.23), reflecting that the non-euro EU members are relatively wealthy economies. When covariate distributions differ substantially between treated and control groups, OLS and gravity models rely on linear extrapolation from regions of the covariate space where control observations are sparse, potentially yielding biased and unstable estimates (Baier and Bergstrand, 2009; Persson, 2001). Baier and Bergstrand (2009) show that matching estimates of FTA treatment effects are much more stable and economically plausible than OLS gravity estimates, which often display extreme instability across years. This motivates our use of causal forests with double machine learning, which extends the matching approach by flexibly adjusting for covariate differences through nonparametric first-stage estimation while also allowing treatment effects to vary with observed characteristics.
Table 4: Covariate Balance: Eurozone vs. Non-Eurozone Pairs (Post-1999)
Variable Eurozone Non-EZ Diff. Std. Diff.
Log bilateral trade 22.27 22.42 -0.15 -0.09
Log GDP product 25.70 26.06 -0.36 -0.23
Log GDP per capita 20.56 20.81 -0.24 -0.46
Pre-euro trade intensity 22.10 22.23 -0.13 -0.09
Note: Standardized difference >> 0.25 indicates meaningful imbalance. Eurozone = pairs where both countries adopted the euro; Non-EZ = pairs involving Denmark, Sweden, or UK. Pre-euro trade intensity is the average log bilateral trade during 1995–1998. Comparison uses post-1999 observations only.
Figures 24 visualize these covariate distributions across key years, following the approach of Baier and Bergstrand (2009). Each panel shows kernel density estimates for eurozone pairs (solid line) and non-eurozone pairs (dashed line). In 1990, before the euro existed, only non-eurozone pairs appear. By 1999, eurozone pairs emerge with distributions largely overlapping non-eurozone pairs across all three covariates. The substantial overlap supports the conditional independence assumption: for most covariate values, we observe both euro and non-euro pairs, enabling credible effect estimation. However, in later years (2005, 2013), non-eurozone pairs show slightly higher GDP and trade levels on average, reflecting that the remaining non-euro EU members (UK, Sweden, Denmark) are relatively large economies. This motivates the flexible nonparametric adjustment provided by the causal forest’s first-stage estimation.
Figure 2: Distribution of Log GDP Product for eurozone pairs (solid) and non-eurozone pairs (dashed) across key years. In 1990, no eurozone pairs exist as the euro had not yet been adopted.
Figure 3: Distribution of Log GDP Per Capita for eurozone pairs (solid) and non-eurozone pairs (dashed) across key years.
Figure 4: Distribution of Log Bilateral Trade for eurozone pairs (solid) and non-eurozone pairs (dashed) across key years.
To further assess overlap between treated and control groups, we estimate propensity scores using both logistic regression and random forest with log GDP product, log GDP per capita, and pre-euro trade intensity as predictors. Figure 5 shows the distribution of propensity scores by treatment status. The logistic regression estimates (left panel) show substantial overlap, while the random forest estimates (right panel) show tighter separation—reflecting RF’s ability to capture nonlinear relationships in treatment assignment. Both treated (eurozone) and control (non-eurozone) pairs span a wide range of propensity scores under both methods, supporting the positivity assumption required for causal inference. The tighter RF distribution is relevant because our DML approach uses flexible ML models in the first stage; the presence of overlap even under RF’s more precise fit is reassuring. Table 5 provides summary statistics.
Figure 5: Propensity score distribution by treatment status. Left panel shows logistic regression estimates; right panel shows random forest estimates. Substantial overlap between treated (eurozone) and control (non-eurozone) pairs supports the positivity assumption.
Table 5: Propensity Score Diagnostics
Statistic Treated Control
N observations 1,155 756
Mean propensity score 0.640 0.550
Std. dev. 0.140 0.139
Min 0.256 0.223
Max 0.950 0.904
Common support region: [0.256, 0.904]
In common support 1,139 (98.6%) 753 (99.6%)
Note: Propensity scores estimated using logistic regression with log GDP product, log GDP per capita, and pre-euro trade intensity as predictors. Treated = eurozone pairs; Control = pairs involving Denmark, Sweden, or UK.
5 Results
5.1 Gravity Benchmark
Before examining heterogeneity, we establish a gravity benchmark using the canonical specification from the euro trade literature. Table 6 presents estimates from several specifications on the same EU15 sample. The two-way fixed effects OLS estimate is 17.0%, while PPML gravity with pair and year fixed effects yields 12.8%. The three-way fixed effects PPML specification—with exporter-time, importer-time, and pair fixed effects following structural gravity best practices (Anderson and van Wincoop, 2003)—yields a higher estimate of 22.6%. These estimates fall within the 4–30% range documented in prior literature (Rose, 2016; Gunnella et al., 2021).
Table 6: Comparison of Euro Trade Effect Estimates
Method Specification Coefficient Effect (%) 95% CI
Two-way FE OLS Pair + Year FE 0.157 +17.0% [0.048, 0.267]
PPML Gravity Pair + Year FE 0.120 +12.8% [0.035, 0.205]
PPML Gravity Three-way FE 0.204 +22.6% [0.123, 0.285]
Causal Forest DML 0.252 +28.6% [0.103, 0.401]
CFFE Node-level FE 0.133 +14.2% [0.103, 0.162]
Notes: All models estimated on EU15 bilateral trade data, 1995–2015. Two-way FE OLS uses pair and year fixed effects with clustered standard errors. PPML = Poisson Pseudo-Maximum Likelihood following Santos Silva & Tenreyro (2006). Three-way FE includes exporter-time, importer-time, and pair fixed effects (Head & Mayer, 2014). Causal Forest uses Double Machine Learning for heterogeneous treatment effects. CFFE = Causal Forests with Fixed Effects. Standard errors clustered at the country-pair level.
The causal forest DML estimate of 29% is somewhat higher than the gravity estimates. This difference reflects two factors. First, the causal forest uses flexible nonparametric first-stage estimation rather than linear fixed effects, potentially capturing nonlinear relationships between confounders and outcomes. Second, and more importantly, the causal forest estimates a different object: while gravity models estimate a single average effect constrained to be constant across pairs, the causal forest estimates heterogeneous effects that can vary with pair characteristics. The ATE from the causal forest is a weighted average of these heterogeneous effects, where the weights depend on the covariate distribution.
The CFFE estimate of 14.1% falls between the gravity and naive causal forest estimates, suggesting that proper fixed effects handling moderates the causal forest estimate toward the gravity benchmark. This convergence is reassuring: when we account for pair and year fixed effects within the causal forest framework, we recover estimates consistent with the gravity literature.
The key advantage of the causal forest approach is not that it produces a different average effect, but that it reveals the distribution of effects underlying that average. The gravity estimate of 10–16% is correct as an average, but it masks substantial heterogeneity that we document below.
5.2 Selection and Identification
As shown in Table 4, treated and control observations differ substantially on observable characteristics. Treated pairs have higher GDP and more trade. This creates selection bias in naive comparisons: eurozone pairs would have traded more than non-eurozone pairs even without the euro, simply because they are larger, richer, and more integrated economies. The direction of bias depends on the estimator. Simple OLS comparisons that fail to control for these differences will overstate the euro effect by attributing pre-existing trade advantages to euro adoption. However, gravity models with extensive fixed effects may understate the effect by absorbing some of the euro’s impact into pair fixed effects, particularly if the euro amplified existing trade relationships (Persson, 2001; Chintrakarn, 2008). The causal forest addresses selection on observables through flexible first-stage estimation that residualizes both the outcome and treatment on confounders, while the heterogeneous effects framework allows us to distinguish between pairs where the euro created new trade versus amplified existing patterns.
5.3 Heterogeneous Treatment Effects
Euro adoption increased bilateral trade by 29% on average (95% CI: $11\%,49\%]$) using the naive causal forest without fixed effects. [^2] Since the outcome variable is log trade, the coefficient estimate is in log points. We convert to percentage change using $(\exp(\hat{\tau})-1)\times 100$. For example, an estimate of 0.252 log points corresponds to $(\exp(0.252)-1)\times 100=29\%$. Figures report estimates in log points; text reports percentage changes. This estimate falls between the gravity model range (4–6%) and the SCM estimates ($\sim$30%) reported by Gunnella et al. (2021). The confidence interval is tighter than in previous studies, reflecting the larger sample and more precise estimation.
The CATE distribution reveals heterogeneity that cannot be captured by methods estimating a single average effect. Among eurozone pairs, effects range from $12\%$ to $79\%$, with substantial variation across pairs. The euro does not have a single effect on trade. Different country pairs experienced dramatically different impacts, from near-zero to substantial gains. Figure 6 shows the distribution for eurozone pairs.
Figure 6: Distribution of Conditional Average Treatment Effects (CATE) for eurozone country pairs. The red dashed line indicates the Average Treatment Effect (ATE). Effects range from $12\%$ to $79\%$ across pairs.
Table 7: Top and Bottom Eurozone Pairs by Euro Trade Effect
Country Pair Effect (%)
Highest Effects
France $\leftrightarrow$ Italy +78.5
Germany $\leftrightarrow$ Italy +78.5
Belgium $\leftrightarrow$ Germany +78.5
Austria $\leftrightarrow$ Germany +78.4
France $\leftrightarrow$ Germany +78.4
Lowest Effects
Greece $\leftrightarrow$ Ireland -9.6
Ireland $\leftrightarrow$ Portugal -10.4
Finland $\leftrightarrow$ Greece -11.4
Greece $\leftrightarrow$ Portugal -12.0
Finland $\leftrightarrow$ Portugal -12.3
Table 7 shows the top and bottom eurozone pairs by estimated euro effect. The highest effects are concentrated among core European pairs (France–Italy at +79%, Germany–Italy at +78%, Belgium–Germany at +78%, Austria–Germany at +78%, France–Germany at +78%). The lowest effects — including some negative point estimates — involve peripheral pairs such as Greece–Ireland ($10\%$) and Ireland–Portugal ($10\%$). Table 8 provides confidence intervals for all pair-level effects. Notably, while some point estimates are negative, most are not statistically distinguishable from zero at the 5% level. However, two pairs—Greece–Portugal and Finland–Portugal—show statistically significant negative effects, suggesting genuine trade diversion for these peripheral pairs.
Table 8: Pair-Level Euro Trade Effects with Confidence Intervals
Country Pair Effect (%) 95% CI Significant
Highest Effects
France $\leftrightarrow$ Italy +78.5% [+60.8%, +98.2%] Yes
Germany $\leftrightarrow$ Italy +78.5% [+60.7%, +98.2%] Yes
Belgium $\leftrightarrow$ France +78.5% [+60.2%, +98.9%] Yes
Belgium $\leftrightarrow$ Germany +78.4% [+60.3%, +98.5%] Yes
Austria $\leftrightarrow$ Germany +78.4% [+60.2%, +98.6%] Yes
France $\leftrightarrow$ Germany +78.4% [+60.6%, +98.1%] Yes
Germany $\leftrightarrow$ Netherlands +78.2% [+60.5%, +97.8%] Yes
Ireland $\leftrightarrow$ Luxembourg +77.5% [+34.2%, +134.6%] Yes
Austria $\leftrightarrow$ Luxembourg +77.1% [+32.4%, +137.0%] Yes
Belgium $\leftrightarrow$ Luxembourg +76.6% [+30.7%, +138.6%] Yes
Lowest Effects
Austria $\leftrightarrow$ Greece -2.2% [-11.0%, +7.5%] No
Austria $\leftrightarrow$ Finland -3.4% [-13.2%, +7.6%] No
Austria $\leftrightarrow$ Ireland -4.6% [-15.4%, +7.7%] No
Austria $\leftrightarrow$ Portugal -6.4% [-15.6%, +3.8%] No
Finland $\leftrightarrow$ Ireland -9.9% [-22.9%, +5.3%] No
Ireland $\leftrightarrow$ Portugal -13.4% [-26.7%, +2.3%] No
Greece $\leftrightarrow$ Ireland -13.4% [-25.8%, +1.0%] No
Finland $\leftrightarrow$ Greece -13.6% [-25.5%, +0.1%] No
Greece $\leftrightarrow$ Portugal -15.1% [-26.6%, -1.9%] Yes (-)
Finland $\leftrightarrow$ Portugal -15.2% [-27.9%, -0.3%] Yes (-)
Notes: Effect shows the average CATE for each country pair. 95% CI from causal forest estimation. Significant indicates whether the 95% CI excludes zero. (+) indicates positive significant effect, (-) indicates negative significant effect.
Table 9 aggregates the pair-level effects to show the average euro effect for each eurozone member country. Luxembourg (+69%), Germany (+50%), Belgium (+47%), and Netherlands (+47%) show the largest average effects, while Greece (+10%) and Finland (+15%) show the smallest. This core-periphery pattern suggests the euro’s benefits were not uniformly distributed.
Table 9: Average Euro Effect by Country (Eurozone Members Only)
Country Effect (%) Std Min (%) Max (%)
Luxembourg +69.0 0.06 +42.7 +85.9
Germany +49.5 0.19 +11.8 +80.3
Belgium +47.3 0.16 -4.9 +85.9
Netherlands +47.3 0.13 +18.9 +82.4
France +36.7 0.20 +8.1 +80.2
Italy +34.8 0.16 +13.2 +78.6
Spain +26.0 0.10 +8.8 +73.3
Austria +22.5 0.24 -17.1 +85.7
Ireland +18.8 0.25 -17.3 +85.7
Portugal +15.8 0.23 -17.4 +71.1
Finland +14.6 0.23 -17.5 +80.3
Greece +9.6 0.21 -17.5 +71.1
Note: Effect shows the average CATE across all eurozone pairs involving each country. Min and Max show the range of pair-level effects. Non-eurozone EU members (Denmark, Sweden, UK) are excluded as they serve as controls.
Figure 7 shows the distribution of effects for each country, revealing substantial within-country heterogeneity. Even countries with high average effects show wide variation across their trading partners.
Figure 7: Distribution of euro trade effects by country. Each box shows the distribution of pair-level CATEs involving that country. The red dashed line indicates the overall ATE.
Figure 8 shows feature importance from the causal forest. Pre-euro trade intensity (63% of feature importance) and combined GDP (27%) are the primary drivers of heterogeneity, together accounting for over 90% of the variation. GDP per capita plays a smaller role (10%).
Figure 8: Feature importance for treatment effect heterogeneity.
To understand how each feature affects treatment effects, Figure 9 shows partial dependence plots for the three main effect modifiers. The relationship between GDP and treatment effects is monotonically positive: larger economies experience larger euro effects. Pre-euro trade intensity shows a similar pattern, with high-trade pairs benefiting most. GDP per capita shows a weaker relationship, consistent with its lower feature importance.
Figure 9: Partial dependence plots showing how each feature affects the predicted treatment effect. Shaded areas indicate 95% confidence intervals. GDP and pre-euro trade intensity show strong positive relationships with treatment effects.
6 Discussion
The heterogeneity we document suggests that the 4–30% range in prior literature appears to reflect genuine variation rather than methodological noise. Previous studies finding widely varying estimates were not producing unreliable results; they were capturing different slices of a heterogeneous distribution. A study focused on core European pairs would find large effects ($\sim$65–68%), while one focused on peripheral pairs would find small or null effects. Our causal forest approach uncovers this full distribution, showing that both findings are correct for their respective subpopulations.
The country-level results reveal a core-periphery pattern. Luxembourg (+69%), Germany (+50%), Belgium (+47%), and Netherlands (+47%) — small, open economies at the geographic heart of the eurozone — benefited most from euro adoption. These countries serve as logistics and financial hubs with trade-to-GDP ratios exceeding 100%, [^3] World Bank data show trade-to-GDP ratios of approximately 400% for Luxembourg, 160% for Belgium, and 150% for the Netherlands in the early 2000s. making them uniquely sensitive to transaction cost reductions. Peripheral economies show smaller effects: Greece (+10%), Portugal (+16%), and Finland (+15%). These countries have smaller domestic markets, fewer natural trading partners within the eurozone core, and greater geographic distance from the European economic center.
Some country pairs show negative point estimates, particularly those involving peripheral countries (Finland–Portugal: $14\%$, Greece–Portugal: $14\%$). While some of these negative estimates are statistically significant, they likely reflect relative trade diversion—peripheral pairs losing market share as firms redirected trade toward core eurozone partners—rather than absolute trade destruction. These pairs also had weak pre-euro trade relationships, leaving little scope for the euro to enhance already-minimal flows.
6.1 Economic Mechanisms: Why Pre-Euro Trade Intensity Matters
The finding that pre-euro trade intensity is the strongest predictor of euro effects raises a natural question: through what economic mechanisms does the euro amplify existing trade relationships rather than create new ones? We consider three complementary explanations.
Following Baldwin (2006), currency unions can expand trade through two channels: the intensive margin (existing exporters ship more) and the extensive margin (new firms begin exporting). Our pair-level patterns are consistent with an intensive-margin-dominant channel. Pairs with high pre-euro trade intensity already had established exporter networks, distribution channels, and customer relationships. For these pairs, the euro reduced transaction costs on existing flows, allowing firms to expand volumes without the fixed costs of market entry. In contrast, pairs with low pre-euro trade lacked these established networks. While the euro reduced variable trade costs, it did not eliminate the fixed costs of entering new markets—learning about foreign regulations, establishing distribution networks, building customer relationships. For peripheral pairs like Finland–Portugal, these fixed costs remained prohibitive even after the euro removed currency friction.
This interpretation aligns with the firm-level evidence in Baldwin et al. (2008), who find that the euro’s trade effects operated primarily through the intensive margin in the early years, with extensive margin effects emerging only gradually. Our pair-level heterogeneity reflects this pattern: core pairs with dense existing trade networks saw immediate intensive-margin gains, while peripheral pairs with sparse networks saw limited benefits because extensive-margin adjustment is slow.
The largest euro effects accrue to pairs embedded in cross-border production networks (Baldwin and Lopez-Gonzalez, 2015). Luxembourg, Belgium, and the Netherlands serve as logistics hubs in European supply chains, with firms conducting frequent small-value transactions across borders. For these transactions, currency conversion costs and exchange rate uncertainty impose disproportionate burdens. Consider a German auto manufacturer sourcing components from Belgian suppliers: before the euro, each shipment faced currency risk; the euro eliminated this friction, enabling tighter supply chain integration and more frequent cross-border transactions.
This mechanism explains why pre-euro trade intensity predicts euro effects: pairs with high pre-euro trade were already integrated into supply chains, and the euro deepened this integration. Pairs with low pre-euro trade were not part of these networks, and the euro alone could not create the complementary investments (logistics infrastructure, supplier relationships, just-in-time systems) needed to join them.
The euro may have generated network effects that reinforced existing trade patterns. As core eurozone pairs deepened their integration, they became more attractive partners for additional trade, potentially diverting trade from peripheral pairs. A French firm choosing between a German and a Portuguese supplier might increasingly favor the German option as euro-denominated supply chains became more efficient. This trade diversion could explain the near-zero or negative effects for peripheral pairs: they lost relative competitiveness as core pairs became more tightly integrated.
Our main finding—that the euro amplified existing trade rather than creating new relationships—appears to conflict with the counterfactual results, where non-adopters’ largest predicted gains are sometimes with weaker trading partners. This apparent contradiction resolves when we distinguish between two types of “weak” partners.
For eurozone members, weak trading partners (e.g., Finland–Portugal) remained weak because the euro could not overcome fundamental barriers: geographic distance, lack of complementary production structures, and absence of established business networks. The euro reduced currency friction but left these deeper barriers intact.
For non-adopters like Sweden and Denmark, the pattern differs because they maintain stable exchange rates with the euro through policy choices (Denmark’s peg) or de facto stability (Sweden’s managed float). Their trade with major eurozone partners (Germany, France) already benefits from low currency friction, so the marginal gain from formal euro adoption is modest. The largest predicted gains come from partners where currency friction remains meaningful—typically smaller eurozone economies where exchange rate management is less precise.
The UK presents a different case: the pound floated freely against the euro, creating genuine currency friction even with major partners like Germany. Euro adoption would have removed this friction across the board, explaining why the UK shows uniformly positive predicted effects with all partners.
The unified interpretation is that the euro’s effect depends on whether it removes a binding constraint. For pairs already enjoying low currency friction (through adoption, pegs, or deep integration), the constraint was never binding, and the euro’s marginal effect is small. For pairs facing genuine currency friction, removing it unlocks trade gains proportional to the underlying economic complementarity between partners.
6.2 Dynamic Heterogeneity: Do Effects Evolve Over Time?
A natural question is whether the heterogeneity we document reflects permanent differences across pairs or differential adjustment speeds. If high-CATE pairs simply adjusted faster to the euro, we would expect the gap between high-CATE and low-CATE pairs to narrow over time as slower-adjusting pairs catch up. Alternatively, if heterogeneity reflects genuine structural differences, the gap should persist.
Table 10 presents CATE estimates separately for three time periods: 1999–2003 (early adoption), 2004–2008 (middle period), and 2009–2015 (late period including the crisis). Figure 10 visualizes these patterns.
Table 10: Dynamic Heterogeneity: Euro Trade Effects by Time Period
Period CATE Group ATE Effect (%) 95% CI N
1999-2003 High-CATE 0.242 +27.4% [0.235, 0.248] 265
Low-CATE 0.218 +24.4% [0.215, 0.221] 260
All pairs 0.230 +25.9% [0.226, 0.234] 525
2004-2008 High-CATE 0.154 +16.6% [0.125, 0.183] 265
Low-CATE 0.230 +25.9% [0.216, 0.245] 260
All pairs 0.192 +21.1% [0.175, 0.208] 525
2009-2015 High-CATE 0.160 +17.3% [0.145, 0.175] 371
Low-CATE 0.189 +20.8% [0.180, 0.198] 364
All pairs 0.174 +19.0% [0.165, 0.183] 735
Notes: Pairs classified as high-CATE or low-CATE based on median split of full-sample predicted treatment effects. Causal forest estimated separately for each time period. Standard errors clustered at pair level. Effect (%) calculated as $(\exp(\text{ATE})-1)\times 100$.
Figure 10: Dynamic heterogeneity: Euro trade effects by time period and predicted CATE group. Pairs classified as high-CATE or low-CATE based on full-sample predicted effects. Error bars show 95% confidence intervals.
The results reveal a nuanced pattern. In the early period (1999–2003), high-CATE pairs show larger effects (+26%) than low-CATE pairs (+20%), consistent with the cross-sectional heterogeneity. However, in the middle period (2004–2008), this pattern reverses: low-CATE pairs show larger effects (+25%) than high-CATE pairs (+14%). By the late period (2009–2015), the original pattern partially re-emerges, with high-CATE pairs at +19% and low-CATE pairs at +15%.
This pattern suggests that the heterogeneity is not simply about adjustment speed. The reversal in the middle period may reflect that low-CATE pairs—which had weaker initial trade relationships—experienced delayed extensive-margin effects as new trade relationships formed. The crisis period (2009–2015) then differentially affected these newer relationships, restoring the original core-periphery pattern. The key finding is that heterogeneity persists across all periods, consistent with genuine structural differences rather than transitory adjustment dynamics.
7 Counterfactual Analysis: Non-Eurozone EU Countries
A key advantage of causal forests is the ability to generate counterfactual predictions for untreated units. We estimate what trade effects Sweden, Denmark, and the United Kingdom would have experienced had they adopted the euro in 1999. Table 11 presents the predicted euro effects for the three EU members that opted out of the eurozone.
Table 11: Predicted Euro Effects for Non-Eurozone EU Countries
Country Effect (%) 95% CI Std Min (%) Max (%)
Sweden +21.9 [+7.7, +38.1] 0.16 -14.9 +85.9
Denmark +19.1 [+4.8, +35.5] 0.19 -16.1 +85.7
United Kingdom +32.6 [+21.4, +44.9] 0.16 +10.3 +80.1
Note: Predicted effects based on causal forest CATE estimates. These are counterfactual predictions for countries that did not adopt the euro. Effect shows the average predicted trade increase if the country had joined the eurozone in 1999. 95% confidence intervals are based on the causal forest variance estimates.
7.1 Counterfactual Support and Validity
A key concern with counterfactual predictions is whether the non-eurozone pairs fall within the support of the training data. If UK, Sweden, or Denmark pairs have characteristics outside the range observed for eurozone pairs, the counterfactual predictions may be unreliable extrapolations. Figure 11 shows the covariate distributions for non-eurozone pairs compared to eurozone pairs used to train the causal forest.
Figure 11: Covariate overlap between non-eurozone pairs (for which we predict counterfactual effects) and eurozone pairs (used to train the model). Substantial overlap in GDP and pre-euro trade distributions supports the validity of counterfactual predictions.
The distributions show substantial overlap, particularly for GDP and GDP per capita. UK pairs tend to have higher GDP product than the average eurozone pair, but remain within the support of the training data. Pre-euro trade intensity shows good overlap for all three countries. Table 12 provides summary statistics on the nearest-neighbor matches, showing that each non-eurozone pair has multiple eurozone pairs with similar characteristics.
Table 12: Counterfactual Support Analysis: Non-Eurozone Country Pairs
Country N Pairs In Support % In Support Avg. Distance
Sweden 11 8 73% 0.38
Denmark 11 5 45% 0.53
United Kingdom 11 11 100% 0.29
Note: “In Support” indicates pairs whose pre-euro covariates (log trade, log GDP product, log GDP per capita) fall within the range observed for eurozone pairs. “Avg. Distance” is the average Euclidean distance (in standardized covariate space) to the nearest eurozone pair. Lower distance indicates better support for counterfactual predictions.
The United Kingdom shows the largest predicted effect (+32.6%, std 0.16, range +10.3% to +80.1%), followed by Sweden (+21.9%, std 0.16, range $14.9\%$ to +85.9%) and Denmark (+19.1%, std 0.19, range $16.1\%$ to +85.7%). These differences reflect each country’s trade structure and existing integration with eurozone partners. Figure 12 summarizes the counterfactual trade trajectories for all three countries, showing actual trade with eurozone partners versus the predicted counterfactual under euro adoption.
Figure 12: Counterfactual trade trajectories for non-eurozone EU members. Each panel shows actual trade with eurozone partners (solid) versus predicted trade under euro adoption (dashed). Trade indexed to 1999 = 100. Shaded areas indicate foregone trade gains.
7.2 Sweden
Sweden’s predicted average effect is +21.9% (std 0.16), with effects ranging from $14.9\%$ to +85.9% across partners. This masks substantial partner-level heterogeneity. Figure 13 shows the predicted effect by trading partner.
Figure 13: Sweden’s predicted euro effect by eurozone trading partner.
Sweden’s largest predicted gains are with Luxembourg (+65%), Austria (+33%), and Belgium (+28%). Sweden-Germany (+13%) and Sweden-Finland (+26%) show moderate predicted effects. Sweden-Greece (+5%) and Sweden-Portugal (+5%) show the smallest effects.
Figure 14 shows the aggregate trade trajectory comparing actual trade to the counterfactual if Sweden had joined in 1999. Figure 15 breaks this down by trading partner, showing actual versus counterfactual trade for each Sweden-eurozone pair.
Figure 14: Sweden-Eurozone trade: actual vs. counterfactual. Left panel shows trade levels; right panel shows cumulative foregone trade.
Figure 15: Sweden’s actual vs. counterfactual trade by eurozone partner. Solid lines show actual trade; dashed lines show predicted trade under euro adoption.
7.3 Denmark
Denmark shows a predicted effect of +19.1% (std 0.19), with effects ranging from $16.1\%$ to +85.7% across partners. Figure 16 shows the partner-level breakdown.
Figure 16: Denmark’s predicted euro effect by eurozone trading partner.
Denmark’s largest gains would have been with Luxembourg (+62%), Belgium (+42%), and Finland (+39%). Denmark-Germany (+11%) shows a smaller effect despite being Denmark’s largest trading partner. Denmark-Portugal ($8\%$) and Denmark-Greece ($3\%$) show negative predicted effects. Denmark’s krone peg to the euro already provides most of the currency stability benefits without full adoption, which may explain the smaller marginal gains with some partners.
Figures 17 and 18 show Denmark’s actual vs. counterfactual trade trajectory at the aggregate and partner levels.
Figure 17: Denmark-Eurozone trade: actual vs. counterfactual. Left panel shows trade levels; right panel shows cumulative foregone trade.
Figure 18: Denmark’s actual vs. counterfactual trade by eurozone partner. Solid lines show actual trade; dashed lines show predicted trade under euro adoption.
7.4 United Kingdom
The United Kingdom shows the largest predicted effect (+32.6%, std 0.16) with uniformly positive effects across all partners, ranging from +10.3% to +80.1%. Figure 19 shows the partner-level breakdown.
Figure 19: United Kingdom’s predicted euro effect by eurozone trading partner.
The UK’s largest predicted gains are with Luxembourg (+71%), Germany (+55%), and Spain (+26%). Even the smallest effects (UK-Belgium at +11%, UK-Ireland at +12%) are positive. UK-eurozone trade appears to have faced significant currency friction that the euro would have removed across the board.
Our UK estimate (+32.6%) is comparable to Saia (2017), who uses the synthetic control method and finds +16%. The difference likely reflects: (1) our method captures partner-level heterogeneity; and (2) SCM constructs a single synthetic counterfactual, while we estimate effects conditional on pair characteristics.
Figures 20 and 21 show the UK’s actual vs. counterfactual trade trajectory at the aggregate and partner levels.
Figure 20: UK-Eurozone trade: actual vs. counterfactual. Left panel shows trade levels; right panel shows cumulative foregone trade.
Figure 21: UK’s actual vs. counterfactual trade by eurozone partner. Solid lines show actual trade; dashed lines show predicted trade under euro adoption.
7.5 Reconciling Amplification vs. Creation Effects
The counterfactual results reveal an interesting pattern. Among eurozone members, pairs with strong pre-euro trade (Belgium-Netherlands, Germany-France) saw the largest effects. Yet for Sweden and Denmark, the largest predicted gains are with weaker trading partners (Italy, France, Spain) rather than stronger ones (Germany, Finland).
The UK presents a different pattern: its largest predicted gains are with Luxembourg (+71%) and Germany (+55%), its largest eurozone partners. This suggests the UK-Germany relationship faced significant currency friction that the euro would have removed.
This apparent contradiction resolves when we distinguish between two mechanisms. Among adopters, the euro amplified trade where integration was already deep — Belgium-Netherlands and Germany-France had extensive supply chains that benefited from eliminating currency risk. For non-adopters, the pattern depends on existing currency arrangements. Sweden and Denmark both maintain relatively stable exchange rates with the euro (Denmark via its peg, Sweden via policy). Their largest partners already enjoy low currency friction, so the marginal gain from euro adoption is smaller. The largest gains come from partners where currency friction remains high. The UK is different: the pound floated freely against the euro, creating currency friction even with major partners like Germany. Euro adoption would have removed this friction across the board.
The unified story is that the euro’s effect is largest where it removes a binding constraint. For already-integrated pairs (whether through adoption or currency pegs), the constraint was never binding. For pairs facing genuine currency friction, removing it unlocks substantial trade gains. Countries with existing currency pegs (like Denmark) may see smaller benefits than those with floating rates.
7.6 Illustrative Counterfactual Exercise
The counterfactual analysis offers suggestive insights for countries considering euro adoption, though these predictions should be interpreted with caution given the strong assumptions required for transportability. The heterogeneity in predicted effects suggests that the decision is not one-size-fits-all.
For the United Kingdom, the predicted +32.6% effect represents the largest potential gain among the three non-eurozone EU members. The uniformly positive effects across all partners (minimum +10.3%) suggest that UK-eurozone trade may have faced substantial currency friction that euro adoption could have reduced. Post-Brexit, this counterfactual becomes moot, but the analysis suggests that the UK’s decision to remain outside the eurozone may have carried some trade costs during its EU membership—though the magnitude is uncertain given the structural differences between the UK and eurozone economies.
For Sweden, the predicted +21.9% effect is substantial but comes with considerable partner-level variation (ranging from $14.9\%$ to +85.9%). Sweden’s largest predicted gains would come from smaller eurozone economies (Luxembourg, Austria, Belgium) rather than its major trading partners. This pattern suggests that Sweden’s existing trade relationships with Germany and Finland may already benefit from low currency friction, potentially limiting the marginal gains from formal euro adoption.
For Denmark, the predicted +19.1% effect is similar to Sweden’s, but the krone’s peg to the euro already captures much of the currency stability benefit. Denmark’s near-zero predicted effects with some peripheral partners suggest that for certain pairs, the benefits of reduced transaction costs may be modest. Denmark’s current arrangement—maintaining the peg without full adoption—may represent a reasonable middle ground, though this is speculative.
These illustrative results suggest that future euro adoption decisions might consider: (1) the extent of existing currency friction with eurozone partners; (2) whether the country’s major trading partners are already in the eurozone; and (3) whether alternative arrangements (like Denmark’s peg) can capture most of the trade benefits without the costs of full monetary union. However, we emphasize that these are suggestive patterns rather than definitive policy recommendations.
Several limitations temper these counterfactual predictions. First, our estimates are backward-looking: they capture the euro’s effects during 1999–2015, a period that included both the pre-crisis boom and the eurozone debt crisis. Future effects may differ as the eurozone’s institutional framework evolves, as new members join, and as global trade patterns shift. The experience of crisis-era adopters (Estonia, Latvia, Lithuania) suggests that adoption timing matters, and countries considering adoption today face a different economic environment than the original 1999 cohort.
Second, our counterfactual analysis assumes that non-eurozone countries would have adopted in 1999 alongside the original members. In practice, later adoption would yield different effects: the eurozone of 2025 differs from the eurozone of 1999 in membership, institutional design, and economic conditions. Countries considering adoption today should not simply extrapolate from our 1999-based counterfactuals.
Third, our partial equilibrium estimates do not account for general equilibrium effects. If the UK had adopted the euro, eurozone trade patterns would have adjusted: some trade currently flowing through non-euro channels might have shifted, and the euro’s overall effect on European trade integration would have differed. Our counterfactual predictions assume the UK’s adoption would not have affected other countries’ trade patterns—an assumption that becomes less tenable for large economies.
Fourth, trade effects are only one consideration in the euro adoption decision. The eurozone crisis demonstrated the costs of losing monetary policy autonomy: countries like Greece, Portugal, and Spain could not devalue their currencies to restore competitiveness, contributing to prolonged recessions. Denmark, Sweden, and the UK maintained independent monetary policy and arguably weathered the crisis better. Our estimates capture trade benefits but not the full cost-benefit calculus of monetary union membership.
Finally, our estimates reflect the euro’s effect on bilateral trade within the EU. They do not capture effects on trade with non-EU partners, foreign direct investment, financial integration, or other economic outcomes that factor into the adoption decision. A comprehensive policy assessment would require integrating our trade estimates with evidence on these other channels.
8 Conclusion
We estimate the full distribution of conditional average treatment effects (CATEs) using causal forests with double machine learning. The approach extends the propensity score matching tradition of Persson (2001) and Chintrakarn (2008) by allowing the effect to vary with observed characteristics, while also offering advantages over synthetic control methods (Saia, 2017; Gunnella et al., 2021) by enabling scalable counterfactual analysis across all country pairs simultaneously.
The results suggest a positive relationship between euro adoption and bilateral trade, but the magnitude varies substantially across country pairs. Our preferred average effect estimate is around 15% (14.1% after fixed effects correction), consistent with gravity benchmarks; the 29% naive estimate reflects heterogeneity-weighted averages that give more weight to high-effect pairs. This heterogeneity explains why prior studies using different samples and methods have produced such divergent estimates. The variation is not methodological noise; it appears to reflect genuine differences in how the euro affected different trading relationships.
Core eurozone pairs with strong pre-existing trade relationships experienced larger effects, while peripheral pairs saw smaller gains. The euro amplified existing trade relationships rather than creating new ones, with pre-euro trade intensity and GDP as key drivers of heterogeneity. Counterfactual analysis suggests non-eurozone EU members would have experienced moderate effects: UK (+32.6%), Sweden (+21.9%), Denmark (+19.1%).
Countries considering euro adoption should expect larger gains if they have strong existing trade ties with core eurozone economies (Germany, France, Netherlands), and smaller gains if they already maintain stable exchange rates with the euro (as Denmark does via its peg). The largest marginal gains come with partners where currency friction currently limits trade expansion. The experience of Greece, Portugal, and Spain during the eurozone crisis illustrates the costs of losing monetary policy autonomy. Denmark, Sweden, and the UK maintained independent monetary policy and arguably weathered the crisis better. Our counterfactual estimates suggest these countries may have forgone some trade gains by staying out, but they retained policy flexibility that proved valuable during economic stress.
Methodologically, this is among the first applications of causal forests to currency union effects, extending the PSM tradition to allow data-driven discovery of effect heterogeneity. The approach also offers advantages over synthetic control methods for counterfactual analysis: whereas SCM requires constructing a separate synthetic counterfactual for each unit of interest, causal forests generate predictions for all untreated units simultaneously once the model is trained. This scalability enabled us to estimate counterfactual effects for three non-eurozone countries across all their eurozone trading partners — an exercise that would require dozens of separate SCM analyses. Empirically, we offer one possible explanation for the 4–30% puzzle: the variation may reflect genuine heterogeneity across country pairs rather than methodological differences alone, though we cannot rule out other explanations. For policy, we provide illustrative counterfactual predictions for non-eurozone EU members with partner-level granularity, subject to the caveats discussed above.
The analysis covers 15 EU countries (EU15) through 2015, matching the methodology of Gunnella et al. (2021). Our estimates suggest that extending to EU28 with naive causal forests produces biased estimates due to crisis-era adoption timing, but the CFFE approach with node-level fixed effects correction recovers estimates consistent with the EU15 baseline. Future work could extend the analysis to include recent euro adopters (Estonia 2011, Latvia 2014, Lithuania 2015) once sufficient post-adoption data accumulates. A causal forest trained on data through 2014 could generate predictions for these recent adopters, enabling out-of-sample validation by comparing predicted effects to actual post-adoption trade changes. The model could also generate predictions for prospective euro members: Romania, Bulgaria, Poland, Czech Republic, Hungary, Sweden, and Denmark. Examining how the euro’s effects changed during the 2008 financial crisis and subsequent debt crisis would shed light on whether the benefits of currency union membership vary with macroeconomic conditions.
Our findings have direct relevance for the seven EU members that have not yet adopted the euro: Bulgaria, Czech Republic, Denmark, Hungary, Poland, Romania, and Sweden. Each faces a distinct cost-benefit calculus shaped by their current trade patterns, exchange rate arrangements, and economic structure. Central European economies (Poland, Czech Republic, Hungary) have deep trade integration with Germany and other core eurozone members, suggesting substantial potential gains from adoption, though these countries also maintain flexible exchange rates that have served as shock absorbers during economic downturns. Southeastern European economies (Bulgaria, Romania) have weaker existing trade ties with the eurozone core, suggesting more modest potential gains. Nordic holdouts (Denmark, Sweden) present interesting cases: Denmark’s euro peg already captures most currency stability benefits, while Sweden’s managed float provides more flexibility but our estimates suggest larger potential gains.
The eurozone of 2025 differs substantially from the eurozone of 1999, with an evolved institutional framework, expanded membership from 11 to 20 countries, and shifted global trade patterns. These changes suggest caution in extrapolating our historical estimates to future adoption decisions. What our analysis does establish is that the euro’s trade effects are heterogeneous, predictable, and substantial for the right country pairs. Countries with strong existing trade ties to the eurozone core, floating exchange rates that create genuine currency friction, and economic structures complementary to European supply chains stand to gain most. The causal forest framework provides a tool for making these assessments with partner-level granularity, moving beyond the single-number estimates that have dominated policy debates.
Our analysis focuses on aggregate bilateral trade flows, but the euro’s effects may vary across product categories. Future research using product-level bilateral trade data could decompose the aggregate effects we document into product-specific components, testing whether the euro disproportionately boosted trade in technology-intensive or differentiated goods.
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
During the preparation of this work the authors used AI-assisted tools (including large language models) in order to assist with code development for the causal forest analysis and data processing, as well as for proofreading, language editing, and reorganizing the manuscript structure. After using these tools, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.
Appendix A.1 Robustness
Our main analysis focuses on the EU15 sample, where treatment timing is balanced and the causal forest produces reliable estimates. We conduct several robustness checks to assess the validity of our identification strategy and the stability of our estimates.
A.1.1 Pre-Trends and Placebo Tests
A key identifying assumption is that treated and control pairs would have followed parallel trends in the absence of treatment. We assess this assumption through event study analysis and placebo tests.
Figure A.1 presents an event study showing estimated effects by year relative to 1999, using 1998 (k=-1) as the reference period. The pre-treatment coefficients (1995–1997) are small in magnitude, ranging from $3\%$ to $5\%$, compared to post-treatment effects that grow to $13\%$ to $24\%$. While a formal joint test of pre-treatment effects rejects the null at conventional levels, the economic magnitude of pre-trends is modest relative to the post-treatment effects. The slight negative pre-treatment coefficients may reflect anticipation effects as countries prepared for euro adoption, or measurement noise in the reference period. Importantly, the pattern shows a clear break at 1999: effects are near zero or slightly negative before adoption, then become positive and grow steadily afterward, consistent with the euro gradually increasing trade as transaction cost reductions compound.
Figure A.1: Event study: estimated euro effects by year relative to 1999. The reference period is 1998 (k=-1). Pre-treatment coefficients (1995–1997) are small in magnitude ($3\%$ to $5\%$) compared to post-treatment effects ($8\%$ to $24\%$). Shaded areas indicate 95% confidence intervals.
We also conduct placebo tests by assigning fake treatment dates before the actual euro adoption. If our estimates reflect genuine euro effects rather than pre-existing trends, we should find no significant “effects” at fake treatment dates. Figure A.2 shows results for placebo treatments in 1995 and 1997. The estimated effects are small and statistically indistinguishable from zero, providing further support for our identification strategy.
Figure A.2: Placebo tests with fake treatment dates. Estimated “effects” for placebo treatments in 1995 and 1997 are not statistically different from zero, supporting the validity of our identification strategy.
A.1.2 Pre-Trends by Predicted CATE Group
A key concern for our heterogeneity analysis is whether high-CATE pairs were already diverging faster than low-CATE pairs before 1999. If so, the heterogeneity we document might reflect pre-existing trends rather than differential euro effects. To address this, we split pairs into top 25% and bottom 25% by predicted CATE and estimate separate event studies for each group.
Table A.1 presents the results. Figure A.3 visualizes the pre-treatment coefficients for both groups.
Table A.1: Pre-Trends Test: Event Study by Predicted CATE Group
High-CATE (Top 25%) Low-CATE (Bottom 25%) Difference
Event Time Coef. SE Coef. SE Diff. p-value
Pre-Treatment Period
k = -4 -0.204 (0.110) 0.048 (0.053) -0.252 0.039
k = -3 -0.192 (0.077) -0.000 (0.036) -0.192 0.024
k = -2 -0.084 (0.093) -0.005 (0.032) -0.079 0.423
k = -1 0.000 (ref) 0.000 (ref)
Post-Treatment Period (selected)
k = +0 -0.157 (0.171) 0.022 (0.031) -0.179 0.304
k = +2 0.070 (0.126) 0.112 (0.045) -0.042 0.757
k = +5 0.288 (0.168) 0.133 (0.085) 0.155 0.409
k = +10 0.197 (0.192) 0.277 (0.116) -0.080 0.720
k = +15 0.171 (0.176) 0.109 (0.116) 0.062 0.769
Joint Test of Parallel Pre-Trends
Wald statistic: 10.03, p-value: 0.018
Result: Pre-trends may differ between groups
Notes: Event study estimates from difference-in-differences regression with pair and year fixed effects, estimated separately for high-CATE (top 25%) and low-CATE (bottom 25%) pairs. Pairs classified by full-sample predicted treatment effects. Reference period is k=-1 (1998). Standard errors clustered at pair level. Joint test examines whether all pre-treatment differences are jointly zero.
Figure A.3: Pre-trends by predicted CATE group. Event study coefficients estimated separately for high-CATE (top 25%) and low-CATE (bottom 25%) pairs. The pre-treatment period (1995–1998) shows that high-CATE pairs had lower trade growth before 1999, suggesting our heterogeneity estimates may be conservative.
The results reveal an interesting pattern. High-CATE pairs show negative pre-treatment coefficients ($8\%$ to $18\%$), while low-CATE pairs show coefficients near zero. A joint test rejects parallel pre-trends at the 5% level (Wald = 10.0, p = 0.018). However, the direction of the difference is the opposite of what would bias our heterogeneity estimates upward: high-CATE pairs were growing slower than low-CATE pairs before 1999, not faster.
This pattern has two implications. First, it suggests our heterogeneity estimates may be conservative: if high-CATE pairs were on a slower trajectory before the euro, the true euro effect for these pairs may be even larger than we estimate. Second, it raises the question of why pairs that would later benefit most from the euro were growing slower beforehand. One interpretation is that these pairs—which tend to be core European pairs with high pre-euro trade intensity—were already near their trade potential under the pre-euro currency regime, leaving less room for growth. The euro then unlocked additional gains by removing the remaining currency friction.
A.1.3 Leave-One-Out Analysis
To assess whether our results are driven by any single country, we conduct leave-one-out analysis, dropping each country in turn and re-estimating the ATE. Figure A.4 shows that the ATE remains stable within the confidence interval of the full-sample estimate regardless of which country is excluded. Notably, dropping Luxembourg—which has the largest estimated effects—reduces the ATE only modestly, and dropping peripheral economies like Greece or Portugal has minimal impact. This stability suggests our findings reflect a robust pattern across the eurozone rather than being driven by outliers.
Figure A.4: Leave-one-out sensitivity analysis. Each point shows the ATE when the indicated country is dropped from the sample. The horizontal line and shaded band indicate the full-sample ATE and 95% CI.
A.1.4 Alternative Outcome Measures
To assess whether our results are sensitive to the choice of outcome variable, we re-estimate the causal forest using alternative measures of bilateral trade. Table A.2 shows results for five outcome specifications. The baseline effect on log bilateral trade is +24.0%. Effects are similar for exports (+17.3%) and imports (+16.7%), with overlapping confidence intervals. This symmetry is reassuring: exports in our data measure the reporter’s outbound trade while imports measure inbound trade, so similar effects suggest the euro boosted trade flows in both directions rather than favoring one side. The slightly larger point estimate for exports is not statistically distinguishable from the imports effect. Year-over-year trade growth shows a small effect (+1.1%), consistent with the level effects accumulating over time. Log trade intensity (trade normalized by GDP) shows a larger effect (+31.1%), suggesting the euro increased trade relative to economic size.
Table A.2: Robustness: Alternative Outcome Measures
Outcome ATE Effect 95% CI N
Log Bilateral Trade (Baseline) 0.215 +24.0% [0.091, 0.339] 1,911
Log Exports Only 0.160 +17.3% [-0.007, 0.326] 3,822
Log Imports Only 0.155 +16.7% [-0.018, 0.327] 3,822
Trade Growth (YoY) 0.011 +1.1 pp [-0.020, 0.043] 1,820
Log Trade Intensity 0.270 +31.1% [0.117, 0.424] 1,911
[…]
A.1.5 Sensitivity to Unobserved Confounding
Following Oster (2019), we assess how much selection on unobservables would be required to explain away our results. Table A.3 presents the analysis. Panel A shows coefficient stability across specifications: the coefficient increases from near-zero without controls to 0.38 with GDP and year fixed effects, then falls to 0.16 with two-way fixed effects. Panel B shows bias-adjusted estimates for different assumptions about the degree of selection on unobservables ($\delta$) and the maximum R-squared ($R_{max}$). The analysis suggests that unobserved confounders would need to be substantially more important than observed confounders to reduce the effect to zero. Since $\delta<0$ (the coefficient increases with controls), omitted variables appear to bias the effect downward rather than upward.
Table A.3: Oster (2019) Sensitivity Analysis
Panel A: Coefficient Stability
Specification Coefficient Effect (%) R2 N
No controls 0.019 +1.9% 0.000 2,149
GDP controls 0.034 +3.5% 0.741 2,149
GDP + Year FE 0.382 +46.5% 0.789 2,149
Two-way FE 0.157 +17.0% 0.603 2,149
Panel B: Bias-Adjusted Estimates ($\beta^{*}$)
$\delta$ $R_{max}$ = 0.8 $R_{max}$ = 0.9 $R_{max}$ = 1.0
0.5 0.385 0.408 0.431
1.0 0.387 0.433 0.479
1.5 0.390 0.459 0.528
2.0 0.392 0.484 0.576
Notes: Panel A shows coefficient stability across specifications. Panel B shows bias-adjusted coefficients ($\beta^{*}$) following Oster (2019). $\delta$ is the ratio of selection on unobservables to observables. $R_{max}$ is the hypothetical R-squared if all relevant variables were included. For $R_{max}=1.0$, $\delta=-0.28$ would be needed to explain away the effect. Since $\delta<0$, the coefficient increases with controls, suggesting omitted variables bias the effect downward.
A.1.6 Time Stability
Table A.4 shows how the EU15 estimate evolves as we extend the sample year by year from 2007 to 2019. The estimates remain remarkably stable, ranging from 22.2% to 28.6% across all time windows. This stability suggests our main findings are not sensitive to the choice of end year and are robust to the inclusion of crisis and post-crisis periods.
Table A.4: Stability of Euro Trade Effect Estimates Across Time Windows (EU15)
Time Period Treated Control ATE 95% CI Effect (%)
1995–2007 572 737 0.244 [0.081, 0.406] 27.6
1995–2008 638 776 0.219 [0.092, 0.346] 24.5
1995–2009 704 815 0.231 [0.076, 0.386] 26.0
1995–2010 770 854 0.200 [0.050, 0.351] 22.2
1995–2011 836 893 0.233 [0.080, 0.385] 26.2
1995–2012 902 932 0.233 [0.084, 0.382] 26.3
1995–2013 968 971 0.237 [0.087, 0.387] 26.7
1995–2014 1,034 1,010 0.238 [0.075, 0.400] 26.9
1995–2015 1,100 1,049 0.252 [0.103, 0.401] 28.6
1995–2016 1,166 1,088 0.228 [0.084, 0.373] 25.7
1995–2017 1,232 1,127 0.250 [0.102, 0.398] 28.4
1995–2018 1,298 1,166 0.232 [0.065, 0.398] 26.0
1995–2019 1,364 1,205 0.227 [0.062, 0.393] 25.5
Notes: Each row shows causal forest estimates using EU15 bilateral trade data for the indicated time period. ATE is the average treatment effect in log points. Effect (%) is $(\exp(\text{ATE})-1)\times 100$.
A.1.7 Extending to EU28: The Puzzle
A natural question is whether our results extend to the full EU28, including countries that joined the EU after 2004 and adopted the euro during the 2008–2015 period. Table A.5 shows how the estimated euro effect changes as we progressively add countries to the EU15 baseline. The results reveal a striking pattern: the ATE remains stable around 31.3–34.1% as we add Slovenia (2007), Cyprus (2008), and Malta (2008), but begins to decline with Slovakia (2009) and drops sharply with the Baltic states (Estonia 2011, Latvia 2014, Lithuania 2015). Adding Lithuania produces an estimate of just 0.9%. By the time we include all EU28 members, the naive causal forest estimate falls to just 4.3% — compared to the EU15 estimate of 28.6%.
Table A.5: Euro Trade Effect: Sensitivity to Sample Composition
Sample N Treated Control ATE Effect (%)
EU15 2,149 1,100 1,049 0.252 28.6
+Slovenia (2007) 2,928 1,520 1,408 0.272 31.3
+Cyprus (2008) 3,272 1,676 1,596 0.293 34.1
+Malta (2008) 3,665 1,844 1,821 0.269 30.9
+Slovakia (2009) 4,087 2,009 2,078 0.256 29.2
+Estonia (2011) 4,498 2,153 2,345 0.181 19.9
+Latvia (2014) 4,954 2,255 2,699 0.165 17.9
+Lithuania (2015) 5,431 2,345 3,086 0.009 0.9
+Croatia (2023) 5,901 2,345 3,556 0.192 21.2
+Poland (never) 6,424 2,345 4,079 0.093 9.7
+Czech Republic (never) 6,936 2,345 4,591 0.003 0.3
+Hungary (never) 7,489 2,345 5,144 -0.070 -6.7
+Romania (never) 8,079 2,345 5,734 0.031 3.2
+Bulgaria (never) 8,654 2,345 6,309 0.042 4.3
Notes: Each row adds one country to the sample. Year in parentheses indicates euro adoption date. Data covers 1995–2019. ATE is the average treatment effect in log points.
This decline is not driven by genuine differences in euro effects across countries. Rather, it reflects a fundamental identification problem: the second-wave adopters joined the eurozone during or immediately after the 2008–2012 financial and sovereign debt crises. Their adoption timing is confounded with adverse macroeconomic conditions that independently depressed trade. The naive causal forest, which does not explicitly control for pair and year fixed effects, attributes some of this crisis-induced trade decline to euro adoption, biasing the estimate downward.
A.1.8 Fixed Effects Correction via CFFE
To address this bias in the EU28 sample, we apply Causal Forests with Fixed Effects (CFFE), following the methodology of Kattenberg et al. (2023). [^4] Kattenberg et al. provide an R package at https://github.com/MCKnaus/causalDML. However, we were unable to compile their package due to missing dependencies in the repository. The first author therefore developed causalfe (Aytuğ, 2026), a Python implementation of CFFE available at https://github.com/haytug/causalfe and installable via pip install causalfe. The key innovation of CFFE is that fixed effects residualization occurs inside each tree node rather than globally before the forest. This node-level approach is theoretically superior because it allows the FE adjustment to adapt to the local covariate distribution within each leaf, avoiding the bias that can arise from global residualization when treatment effects are heterogeneous. CFFE residualizes both the outcome (log trade) and treatment (euro adoption) on pair and year fixed effects within each tree node before estimating the local treatment effect. This removes time-invariant pair characteristics and common year shocks while preserving the ability to estimate heterogeneous effects.
CFFE produces substantially lower estimates than the naive causal forest for both samples: 14.1% vs 29% for EU15. This difference reflects the removal of pair-specific heterogeneity that can inflate naive estimates when high-trade pairs (which tend to have larger effects) are overweighted. The key distinction between samples is the direction of bias in naive estimates. For EU15, where all eurozone members adopted in 1999–2001 before major crises, the naive estimate is biased upward. For EU28, which includes countries that adopted during the 2008–2012 crisis period, the naive estimate is biased downward because crisis-induced trade declines are conflated with treatment effects.
Table A.6 shows the results. The naive EU28 estimate of 4.3% has a wide confidence interval due to crisis-era adopters. After CFFE correction, the EU28 estimate is 13.4% [12.1%, 14.8%] — close to the EU15 CFFE baseline of 14.1%.
Table A.6: Euro Trade Effect Estimates: Method Comparison (EU28)
Method ATE Effect (%) 95% CI
Naive CF (EconML) 0.042 4.3% [-52.1%, 127.3%]
CFFE (node-level FE) 0.126 13.4% [12.1%, 14.8%]
Notes: Naive CF = Causal Forest (EconML) results from Table A.5. CFFE = Causal Forests with Fixed Effects, using node-level two-way FE residualization. ATE is the average treatment effect in log points. Effect (%) is $(\exp(\text{ATE})-1)\times 100$. Standard errors for CFFE are cluster-robust at the pair level. Sample: EU28 countries, 1995–2019.
This convergence is reassuring. It suggests that the true euro effect is similar across EU15 and EU28 countries, and the apparent decline in the naive EU28 estimate was indeed driven by confounding from crisis-era adoption timing rather than genuine differences in euro effects. After controlling for fixed effects, the euro’s trade-creating effect is approximately 13–14% regardless of sample composition. These results further support our interpretation that average estimates are highly sensitive to the composition of treated units, reinforcing the importance of modeling heterogeneous treatment effects.
A.1.9 Heterogeneity in the EU28 Sample
The CFFE estimates for EU28 also reveal substantial heterogeneity, consistent with our EU15 findings. Figure A.5 shows the distribution of CATEs across all observations. The distribution shows meaningful variation in treatment effects across country pairs, with the interquartile range spanning from modest to substantial gains.
Figure A.5: Distribution of Conditional Average Treatment Effects (CATE) for EU28 sample using CFFE. The red dashed line indicates the ATE.
Figure A.6 shows the distribution of effects by country. The pattern differs from EU15, with newer eurozone members showing the largest effects: Malta (+41.2%), Cyprus (+32.0%), and Estonia (+26.4%) lead the rankings, while original eurozone members Portugal (+10.7%) and Greece (+10.8%) show the smallest effects. The overall ATE is +13.4%. Countries that adopted during the crisis period show smaller gains, aligning with economic intuition: countries that adopted during adverse macroeconomic conditions and had weaker pre-existing trade ties with the eurozone core experienced smaller benefits. Table A.7 provides the full country-level breakdown.
Figure A.6: Distribution of euro trade effects by country (EU28 CFFE). Each box shows the distribution of pair-level CATEs involving that country. The red dashed line indicates the overall ATE.
Table A.7: Euro Trade Effect by Country (EU28 CFFE)
Country Effect (%) N Pairs
Malta +41.2 18
Cyprus +32.0 18
Estonia +26.4 18
Luxembourg +23.6 18
Latvia +22.0 18
Slovenia +21.0 18
Lithuania +17.6 18
Germany +15.3 18
France +14.5 18
Italy +13.8 18
Netherlands +13.4 18
Spain +12.9 18
Slovakia +12.8 18
Belgium +12.3 18
Ireland +12.2 18
Austria +12.0 18
Finland +11.9 18
Greece +10.8 18
Portugal +10.7 18
Overall ATE +13.4
A.1.10 Internet Adoption as a Potential Confounder
A potential concern is that internet adoption followed a similar trajectory to euro adoption: negligible before 1995, then growing rapidly in a convex pattern through the early 2000s. If internet adoption independently boosted trade, and its timing resembles the euro treatment, our estimates could be confounded.
We address this concern using World Bank data on internet usage (individuals using the internet as a percentage of population) from 1990–2015. Figure A.7 shows that internet adoption does follow a convex trajectory around 1999, superficially resembling the euro treatment timing.
Figure A.7: Internet adoption trajectory vs euro treatment. Panel A shows internet penetration by country (blue = eurozone, red dashed = non-eurozone). Panel B compares the normalized internet trajectory with the euro dummy (0/1). While both show increases around 1999, the patterns differ.
However, if internet adoption were driving our heterogeneous euro effects, we would expect countries with higher internet penetration to show larger CATEs. Table A.8 tests this prediction by correlating country-level average CATEs with internet adoption measures across different time periods.
Table A.8: Euro Effect vs Internet Adoption: Eurozone Countries
Country Avg CATE Internet 1999-2002 (%) Internet 2000-2015 (%)
Spain 0.181 14.8 52.5
Portugal 0.165 17.1 42.9
Italy 0.164 23.2 42.7
France 0.163 20.0 57.8
Germany 0.158 32.9 69.2
Greece 0.151 10.4 36.7
Netherlands 0.142 48.5 79.2
Belgium 0.133 30.2 64.4
Finland 0.121 43.8 76.0
Austria 0.116 33.1 64.0
Ireland 0.111 19.4 56.0
Luxembourg 0.099 29.1 73.0
Correlation with CATE:
     Internet 1999–2002 (early adoption): r=0.42r=-0.42
     Internet 2000–2015 (full post-treatment): r=0.53r=-0.53
Note: This table tests whether the estimated euro effects could be confounded by internet adoption. The early-period correlation (1999–2002) tests whether countries with higher internet penetration at the time of euro adoption showed larger effects; this correlation is weak. The full post-treatment correlation is stronger, but this likely reflects reverse causality: countries that benefited more from the euro also developed faster economically, leading to higher internet adoption. Crucially, Luxembourg has the highest CATE but not the highest internet penetration in either period, while Finland and Netherlands have high internet but only average CATEs.
The appropriate test for confounding is whether countries with higher internet penetration at the time of euro adoption showed larger effects. Among eurozone members, the correlation between CATEs and internet penetration during the early adoption period (1999–2002) is actually negative (r=0.42r=-0.42). The full post-treatment period (2000–2015) shows a similar negative correlation (r=0.53r=-0.53). These negative correlations indicate that countries with higher internet penetration tended to have smaller euro effects, the opposite of what we would expect if internet were driving the results.
The pattern of heterogeneity is inconsistent with internet driving the results. Spain has the highest CATE (0.18) but only moderate internet penetration—15% in 1999–2002 and 53% in 2000–2015. In contrast, Netherlands has the highest early internet penetration (48%) but only an average CATE (0.14). If internet were driving the results, we would expect these rankings to align; they do not.
Figure A.8 visualizes these correlations. Panel A shows the negative early-period correlation (r=0.42r=-0.42), while Panel B shows the negative full post-treatment correlation (r=0.53r=-0.53). In both panels, countries with high internet penetration (Netherlands, Finland) appear near the bottom of the CATE distribution, while countries with lower internet penetration (Spain, Portugal) show higher CATEs.
Figure A.8: Euro effect (CATE) vs internet adoption. Panel A plots average CATE against internet penetration during the early adoption period (1999–2002), showing a negative correlation (r=0.42r=-0.42). Panel B plots CATE against internet penetration over the full post-treatment period (2000–2015), showing a similar negative correlation (r=0.53r=-0.53). Countries with high internet penetration (Netherlands, Finland) have lower CATEs, while countries with lower internet (Spain, Portugal) have higher CATEs.
This evidence suggests that internet adoption is not confounding our results. The negative correlations and the misalignment between internet and CATE rankings indicate that the euro effects we estimate reflect genuine currency union effects rather than confounding from concurrent technological change.
The literature on internet and trade finds modest effects. Freund and Weinhold (2004) estimate that a 10 percentage point increase in internet growth raises export growth by 0.2 percentage points. More recent work using 1996–2014 data finds elasticities of 0.03–0.13%, with larger effects among high-income countries (Rodriguez-Crespo et al., 2021). Even at the upper bound, these magnitudes are too small to explain our estimated euro effects of 12–48%. The negative correlation between internet and CATEs (r=0.53r=-0.53) further suggests that internet is not driving our results.
As a more direct test, we re-estimate our causal forest model adding pair-level internet penetration (the average of both countries’ internet usage) as a fourth effect modifier alongside GDP, GDP per capita, and pre-euro trade intensity. If internet adoption were confounding our estimates, controlling for it should substantially reduce the estimated euro effect. Instead, the ATE changes minimally: from 0.252 (28.6%) in the original specification to 0.255 (29.0%) with internet included—a change of only 1.2%. The confidence intervals overlap almost entirely ([0.103, 0.401] vs [0.101, 0.409]). Moreover, the individual CATEs from both models are highly correlated ($r=0.94$), indicating that the heterogeneity patterns are robust to controlling for internet. Table A.9 presents the full comparison, and Figure A.9 visualizes these results.
We also conduct this robustness check using our CFFE estimator on the EU28 sample. Here, adding internet as a fourth covariate reduces the ATE from 0.126 (13.4%) to 0.100 (10.5%)—a more substantial 20% reduction. This larger sensitivity in the CFFE specification likely reflects the expanded EU28 sample, which includes Eastern European countries where internet adoption and EU/euro accession occurred more simultaneously. The confidence intervals still overlap ([0.114, 0.138] vs [0.088, 0.112]), and the effect remains statistically significant and economically meaningful. Importantly, even after controlling for internet, the euro effect remains positive and significant at approximately 10.5%, consistent with genuine currency union benefits. Table A.10 presents the CFFE comparison.
Table A.9: Causal Forest Estimates With and Without Internet as Covariate
Original With Internet
(3 covariates) (4 covariates)
Average Treatment Effect
ATE (log points) 0.252 0.255
95% CI [0.103, 0.401] [0.101, 0.409]
Effect (%) 28.6% 29.0%
CATE Distribution
Mean 0.252 0.255
Std. Dev. 0.206 0.205
Min -0.192 -0.164
Max 0.620 0.648
Feature Importance
Log GDP 0.221 0.191
Log GDP per capita 0.050 0.027
Pre-euro trade 0.729 0.713
Internet penetration 0.069
Model Comparison
CATE correlation 0.96
ATE change $1.2\%$
Notes: Original model uses GDP, GDP per capita, and pre-euro trade intensity as effect modifiers. The augmented model adds pair-level internet penetration (average of both countries’ internet usage). The high CATE correlation and minimal ATE change indicate that internet adoption does not confound the estimated euro effects. Sample: EU15 countries, 1995–2015.
Table A.10: CFFE Estimates With and Without Internet as Covariate (EU28)
Original CFFE With Internet
(3 covariates) (4 covariates)
Average Treatment Effect
ATE (log points) 0.130 0.093
Std. Error 0.006 0.006
95% CI [0.118, 0.142] [0.081, 0.105]
Effect (%) 13.9% 9.8%
Model Comparison
ATE change $28.3\%$
N observations 8,694
Notes: CFFE estimates on EU28 sample. Original model uses GDP, GDP per capita, and pre-euro trade intensity as effect modifiers. The augmented model adds pair-level internet penetration (average of both countries’ internet usage). The larger sensitivity compared to the EU15 causal forest results likely reflects the inclusion of Eastern European countries where internet adoption and EU/euro accession occurred more simultaneously. Despite the reduction, the euro effect remains positive and statistically significant.
Figure A.9: Comparison of causal forest estimates with and without internet as a covariate. Panel A shows the ATE is virtually unchanged (28.6% vs 29.0%). Panel B shows feature importance: internet has high importance but does not displace the other covariates. Panel C shows CATEs from both models are highly correlated ($r=0.94$), confirming that controlling for internet does not alter the heterogeneity patterns.
A.1.11 Trade Diversion
A natural concern is whether the positive intra-eurozone trade effects come at the expense of trade with non-eurozone partners—the classic trade diversion concern from customs union theory. If the euro simply redirected trade from non-eurozone to eurozone partners, the welfare implications would be less favorable than if it created genuinely new trade.
We test for trade diversion by estimating the euro’s effect on eurozone countries’ trade with non-eurozone EU partners. Using our EU15 data, we examine whether eurozone countries reduced trade with the UK, Sweden, and Denmark after adopting the euro.
Table A.11 presents the results. The intra-eurozone effect (+35.5%) represents trade creation—the positive effect of both countries sharing the euro. For trade from eurozone to non-eurozone EU members, we find a positive but imprecisely estimated effect (+5.5%, 95% CI: -63.1% to +201.7%). For trade from non-eurozone EU members to eurozone countries, the effect is also positive (+34.9%, 95% CI: -2.2% to +86.1%).
Table A.11: Trade Creation vs Trade Diversion
Pair Type Effect 95% CI N Obs N Pairs
Intra-Eurozone +35.5% [-0.4%, +84.4%] 1,155 55
Eurozone → Non-EZ EU +5.5% [-63.1%, +201.7%] 504 24
Non-EZ EU → Eurozone +34.9% [-2.2%, +86.1%] 189 9
Notes: Intra-Eurozone measures the effect of both countries adopting the euro (trade creation). Eurozone → Non-EZ EU measures the effect of the reporter country adopting the euro on trade with non-eurozone EU members (UK, Sweden, Denmark). Non-EZ EU → Eurozone measures the reverse direction. A negative effect in the cross-eurozone pairs would indicate trade diversion. Estimated using CausalForestDML on EU15 data, 1995–2015.
The wide confidence intervals for cross-eurozone trade reflect limited statistical power: the number of non-eurozone EU partners (UK, Sweden, Denmark) is small. However, the point estimates are uniformly positive, providing no evidence of trade diversion. If anything, eurozone countries appear to have maintained or increased trade with non-eurozone EU partners after euro adoption, though only the intra-eurozone effect is precisely estimated.
This pattern is consistent with the euro reducing transaction costs for all trade involving eurozone countries, not just intra-eurozone trade. A German firm that adopts euro-denominated invoicing for French customers may also find it easier to quote prices to British or Swedish customers in a stable currency. The absence of trade diversion suggests the euro’s trade effects are primarily trade-creating rather than trade-diverting, though we cannot rule out small diversion effects given the imprecision of our cross-eurozone estimates.
Appendix A.2 Computational Details
This appendix provides technical details on the causal forest implementation, including runtime, convergence diagnostics, and sensitivity to random seeds.
A.2.1 Implementation
All analyses were conducted in Python 3.10 using the EconML library (version 0.14.1) for CausalForestDML estimation. The CFFE analysis uses the causalfe package (version 0.2.0), a Python implementation of Causal Forests with Fixed Effects developed by the first author and available at https://github.com/haytug/causalfe.
A.2.2 Hyperparameters
The CausalForestDML estimator was configured with the following hyperparameters:
  • 1.
    First-stage outcome model: Random Forest with 200 trees, min_samples_leaf = 20
  • 2.
    First-stage treatment model: Random Forest Classifier with 200 trees, min_samples_leaf = 20
  • 3.
    Causal forest: 500 trees, min_samples_leaf = 30, honest splitting enabled
  • 4.
    Cross-fitting: 5-fold cross-validation for nuisance estimation
The CFFE estimator used similar settings with the addition of node-level fixed effects residualization for pair and year effects.
A.2.3 Runtime
Table A.12 reports computation times for the main analyses on a standard workstation (Apple M1 Pro, 16GB RAM). The EU15 causal forest completes in under 30 seconds, while the EU28 CFFE analysis requires approximately 2 minutes due to the larger sample and fixed effects computation.
Table A.12: Computation Times
Analysis Sample Size Runtime (seconds)
EU15 Causal Forest 2,189 28
EU15 CFFE 2,189 45
EU28 Causal Forest 8,456 52
EU28 CFFE 8,456 124
Counterfactual predictions 3
Leave-one-out (14 iterations) 392
A.2.4 Convergence
We assess convergence by examining how estimates stabilize as the number of trees increases. The ATE estimate stabilizes after approximately 200 trees, with minimal variation beyond 300 trees. Our choice of 500 trees provides a comfortable margin for convergence.
The confidence interval width also stabilizes with increasing trees, declining from approximately 0.45 log points with 50 trees to 0.33 log points with 500 trees. Additional trees beyond 500 provide diminishing returns in precision.
A.2.5 Seed Sensitivity
To assess sensitivity to random initialization, we re-estimated the EU15 causal forest with 20 different random seeds. Table A.13 reports the distribution of ATE estimates across seeds.
Table A.13: Seed Sensitivity Analysis (20 seeds)
Statistic Value
Mean ATE 0.207
Std. Dev. 0.008
Minimum 0.193
Maximum 0.221
Coefficient of Variation 3.9%
The ATE estimates are highly stable across seeds, with a coefficient of variation of only 3.9%. The range of estimates (0.193 to 0.221) falls well within the confidence interval of any individual estimate, indicating that our findings are not sensitive to random initialization.
A.2.6 Reproducibility
All code and data necessary to reproduce the analyses are available at [repository URL to be added upon publication]. The random seed for the main results reported in the paper is 42. Running the analysis with this seed will reproduce the exact estimates reported in the tables and figures.